OpenStudy (anonymous):

Write an equation in slope-intercept form for the line with slope 1/2 that contains the poiint (4,6)

OpenStudy (anonymous):

first the slope and a point equation y-y1=m(x-x1)

OpenStudy (anonymous):

y-6=1/2(x-4)

OpenStudy (anonymous):

y-6=1/2x-2

OpenStudy (anonymous):

y=1/2x+4

OpenStudy (anonymous):

Is that the answer and thanks you very much

OpenStudy (anonymous):

u r welcome